# Extracting the expression for the field equations of a diffeomorphism invariant theory of gravity from surface term

### v1

@@ 7 The procedure for extracting the expression of the field equations out of the surface term
For any scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$, where the nonnegative integer $m$ and each of the tensors $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s are allowed to be arbitrary, the surface term $\Theta^{\mu}_{gen}$ arising from such a scalar is presented by equation (86), which can be always decomposed into the form (92) in terms of the second-rank tensors $\check{X}^{\mu\nu}$ and $\check{Q}^{\mu\nu}$ under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$. Particularly, when the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is embedded in equation (83) representing the variation of the Lagrangian density ${L}_{gen}$ within equation (82), the $\check{X}^{\mu\nu}$ tensor resulting from the surface term must be able to participate in the construction of the expression ${E}^{\mu\nu}_{gen}$ for the field equations corresponding to the Lagrangian $\sqrt{-g}{L}_{gen}$ in such a way that ${E}^{\mu\nu}_{gen}=\check{X}^{\mu\nu}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}$ under the condition that equation (93), i.e. equation (83) with $\delta$ substituted by $\mathcal{L}_{\zeta}$, is valid.
@@ A.1 General formalism
In accordance with the spirit of the work [20], here we stress that the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$ performed towards the variation equation (97) of the Lagrangian is supposed to be allowed so that Eq. (100) holds identically throughout the present appendix. Owing to the disappearance of the right hand side of Eq. (100), then utilizing Poincare lemma gives rise to
@@ A.2 The application in Lagrangian (1)
$\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ $=$ $-\big({B}^{\mu\nu}-{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}\big)\mathcal{L}_{\zeta}{g}_{\mu\nu}$
$+2\nabla_{\mu}\big[\zeta_{\nu}\big({P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}+{W}^{\mu\nu}\big)\big]\,$,
@@
Actually, on the basis of the fact that the scalar $\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ originates from the variation of the Lagrangian density L{L}, by utilizing the identity (LABEL:SumZiLieR), we can offer another explanation for that the coexistence of the two equations (39) and (52) is a sufficient condition for giving rise to Eq. (25) as shown by Eq. (64). For the sake of doing so, direct computation shows that the scalar $\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is associated with the covariant divergence of the surface term $\Theta^{\mu}$ in the manner

### v2

@@ 7 The procedure for extracting the expression of the field equations out of the surface term
For any scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$, where the nonnegative integer $m$ and each of the tensors $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s are allowed to be arbitrary, the surface term $\Theta^{\mu}_{gen}$ arising from such a scalar is presented by equation (86), which can be always decomposed into the form (93) in terms of the second-rank tensor $\check{X}^{\mu\nu}$ and the anti-symmetric one $\check{Q}^{\mu\nu}$ under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$. Particularly, when the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is embedded in equation (83) representing the variation of the Lagrangian density ${L}_{gen}$ within equation (82), the $\check{X}^{\mu\nu}$ tensor resulting from the surface term must be able to participate in the construction of the expression ${E}^{\mu\nu}_{gen}$ for the field equations corresponding to the Lagrangian $\sqrt{-g}{L}_{gen}$ in such a way that ${E}^{\mu\nu}_{gen}=\check{X}^{\mu\nu}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}$ under the condition that equation (95), i.e. equation (83) with $\delta$ substituted by $\mathcal{L}_{\zeta}$, is valid. Besides, the anti-symmetric tensor $\check{Q}^{\mu\nu}$ can be regarded as the Noether charge two-form associated to this Lagrangian.
@@ A.1 General formalism
In accordance with the spirit of the work [20], here we stress that the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$ performed towards the variation equation (100) of the Lagrangian is supposed to be allowed so that Eq. (103) holds identically throughout the present appendix. The identity (103) is instrumental in deriving the main results in this appendix. Owing to the disappearance of the right hand side of Eq. (103), then utilizing Poincare lemma gives rise to
@@ A.2 The application in Lagrangian (1)
$\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}=$
$-\big({B}^{\mu\nu}-{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}\big)\mathcal{L}_{\zeta}{g}_{\mu\nu}$
$+2\nabla_{\mu}\big[\zeta_{\nu}\big({P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}+{W}^{\mu\nu}\big)\big]\,$, (133)
@@
Actually, on the basis of the fact that the scalar $\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ originates from the variation of the Lagrangian density L{L}, by utilizing the identity (133), we can offer another explanation for that the coexistence of the two equations (39) and (52) is a sufficient condition for giving rise to Eq. (25) as shown by Eq. (64). For the sake of doing so, direct computation shows that the scalar $\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is associated with the covariant divergence of the surface term $\Theta^{\mu}$ in the manner